The derivate of inv function of f:|0,1 |➡️R f(x) =xe^x at x=0.5 is


1
Expert's answer
2021-02-24T07:09:21-0500

"f(x)=x*e^x"

"f'(x)=e^x+x*e^x"

"f^{-1}(x) -\text{inverse function }f(x)"

"(f^{-1}(x))'=\frac{1}{f'(f^{-1}(x))}"

"f^{-1}(x):x=y*e^y"

"x=y*e^y - \text{this function does not allow you to explicitly express y in x}"

"x\approx{y*(1+y) }"

"y\approx\frac{-1+\sqrt{1+4*x}}{2}"

"f^{-1}(0.5)\approx0.37"

"(f^{-1}(0.5))'=\frac{1}{f'(0.37)}=\frac{1}{e^{0.37}+0.37*e^{0.37}}\approx 0.504"

Answer:0.504